OpenStudy (anonymous):

-2 X -3/4 X 1/6 A. -1/4 B. -1/6 C. 1/6 D. 1/4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

\[-2*\frac{ -3 }{ 4 }*\frac{ 1 }{ 6 }= \frac{ -3 }{ 2 }*\frac{ 1 }{ 6 }=\frac{ -1 }{ 4 }\]

OpenStudy (anonymous):

sorry 1/4

OpenStudy (anonymous):

ok thank you for the help :)

OpenStudy (anonymous):

:)